# NestedFP: High-Performance, Memory-Efficient Dual-Precision Floating Point Support for LLMs

### v1

@@ 4.3 GPU Kernel Design
While executing FP8 GEMM in NestedFP is straightforward—simply performing GEMM on the NestedFP upper tensor, FP16 GEMM is more complex. Figure 6 illustrates the FP16 GEMM process in both a standard FP16 CUTLASS kernel and the NestedFP16 kernel. In the standard FP16 kernel, weights are loaded into shared memory, transferred to registers, and directly consumed by tensor cores without any transformation. In contrast, the FP16 path in NestedFP requires a reconstruction step before tensor core execution; weights are stored as two 8-bit tensors, which must be loaded into shared memory, copied to registers, and then reconstructed into full FP16 values through SIMT-level bitwise operations. Only after this reconstruction can the data be used by the tensor cores.
This additional step introduces a source of overhead absent in traditional FP16 pipelines. To prevent it from becoming a performance bottleneck, SIMT operations must be highly efficient, and their latency must be effectively hidden within the broader compute pipeline. We present two key optimization strategies to address this challenge: (1) SIMT operation optimization and (2) pipelining and scheduling optimizations to prevent SIMT operations from stalling the tensor core pipeline.
SIMT Reconstruction Operation Optimization. To mitigate the overhead introduced by SIMT operations, we fuse four 8-bit bitwise operations into a single 32-bit operation, following a strategy similar to any_precision_llm ; fp6_llm . As shown in Figure 6, the reconstruction process combines two 8-bit tensors, $W_{1}$ and $W_{2}$, stored in registers, into a full 16-bit FP16 weight. It begins with a correction step for $W_{1}$, followed by dropped $E_{1}$ restoration, and ends with the final fusion of $W_{1}$ and $W_{2}$. In this scheme, the lower 8 bits of the FP16 mantissa, $M[3:10]$, are taken directly from $W_{2}$. The upper tensor, $W_{1}$, contains $E[2:5]$ and $M[1:2]$, which may require correction depending on rounding. Rounding is detected using a checksum condition: the LSB of $W_{1}$ ($M_{3}^{\prime}$) and the MSB of $W_{2}$ ($M_{3}$) should match. If $M_{3}^{\prime}=1$ and $M_{3}=0$, no correction is needed for $E[2:5]$ and $M[1:2]$. If $M_{3}^{\prime}=0$ and $M_{3}=1$, rounding may have corrupted $E[2:5]$ and $M[1:2]$, and correction is applied by subtracting 1 at the $M_{3}$ bit position. To unify both cases and eliminate branching, we apply a branch-free correction; by subtracting $M_{3}$ from $W_{1}$ and taking only bits that corresponds to bit position $E[2:5]$ and $M[1:2]$ from $W_{1}$. Omitted zero is then restored and the full FP16 value is reconstructed using NVIDIA’s __byte_perm instruction.
@@
Figure 6: Algorithm of SIMT-Efficient Bitwise Reconstruction of FP16 Weights.
Pipelining and Scheduling Optimizations. The standard FP16 GEMM kernel sources weight elements from registers and employs a two-stage pipeline, overlapping shared memory-to-register transfers with tensor core execution. Following prior work that integrates SIMT operations into the tensor core pipeline fp6_llm , we extend this design to a three-stage pipeline that concurrently overlaps register to shared memory transfers, SIMT operations, and tensor core execution. However, simply adopting a three stage pipeline is insufficient to fully hide the overhead of SIMT operations. Through empirical analysis, we find that distinct optimization strategies are required for non-cooperative and cooperative kernels. In the non-cooperative kernel, four warps jointly execute a single warp group MMA (Matrix Multiply Accumulate), whereas the cooperative kernel employs eight warps to execute two warp group MMAs concurrently.
While shared memory transfers and tensor core executions are asynchronous operations—allowing subsequent instructions to be issued without waiting for completion—SIMT operations are synchronous and block instruction issue during their execution. In the non-cooperative kernel, we advance the scheduling of shared memory to register transfers relative to the naive three stage pipeline and issue multiple copies in bulk. This approach not only allows multiple shared memory to register copy concurrent with SIMT execution, but also allows SIMT instruction to be issued without delay as operands are pre-loaded into registers. However, in the cooperative kernel, which leverages additional compute parallelism by executing two warp group MMAs concurrently, warp groups contend for shared resources such as SIMT pipelines, tensor core pipelines, and shared memory bandwidth. This contention reduces the effectiveness of the optimizations used in the non-cooperative kernel. To address this, we explored several alternatives and found that inserting a compiler-level fence—specifically, an NVVM code motion barrier—after the transformation stage significantly improves performance. This fence inhibits compiler optimizations that may delay or reorder memory operations and serves as a signal for the imminent use of a transformed input in a register, preventing the compiler from eliminating or deferring its materialization.
@@ 5 Evaluation
5.1 NestedFP8 Accuracy
@@
To evaluate the accuracy of our NestedFP8 model, we compare it with an FP8 model using commonly adopted quantization configurations—per-channel weight quantization and per-token activation quantization, both scaled by the absolute maximum value. We evaluate on four models (Llama 3.1 8B, Mistral Nemo (12B), Phi-4 (14B), and Mistral Small(24B)) using three downstream tasks: Minerva Math, MMLU Pro, and BBH. All models are integrated into the vLLM framework, and accuracy is assessed using the LM Evaluation Harness lm-evaluation-harness framework. As noted in Section 4.2, 8.75% of linear layers in Phi-4 are not applicable to NestedFP and are thus executed in FP16. Our method internally applies a single global scaling factor (256) to all weight tensors. For activation quantization, we apply per-tensor scaling based on the absolute maximum value. As shown in Table 2, our method achieves accuracy comparable to the FP8 baseline despite foregoing fine-grained quantization, demonstrating that NestedFP8 can preserve accuracy.
@@
Llama 3.1 8B Mistral Nemo Phi-4 Mistral Small FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) Minerva Math 17.6 17.2 16.9 16.8 42.9 41.3 34.6 35.7 MMLU Pro 32.8 33.0 34.9 35.0 52.7 53.3 53.6 53.7 BBH 38.6 38.8 41.2 40.6 27.1 27.9 51.8 51.4
Table 2: Accuracy comparison between FP8 baseline and NestedFP8. FP8(B) denotes FP8 baseline, and FP8(N) denotes NestedFP8.
5.2 NestedFP16 Kernel Performance
Comparison over Vanilla FP16 Kernel. To compare the performance of NestedFP16 kernel with normal FP16 GEMM kernel fairly, we implemented a highly optimized FP16 baseline using the same CUTLASS codebase. This baseline selects the best configuration for each GEMM shape through exhaustive search over the design space, closely matching cuBLAS performance (validated in Appendix). Our kernel design space includes non-cooperative and cooperative variants. For non-cooperative kernels, we perform a grid search over tile shapes with $T_{m}\in{16,32,64,128,256}$, $T_{n}\in{64,128,256}$, and $T_{k}\in{64,128,256}$. Cooperative kernels use larger tiles ($T_{n}\in{128,256}$), and configurations that fail to compile are excluded. Non-cooperative kernels use $(1,1,1)$ clusters, while cooperative ones use $(2,1,1)$, assigning two thread blocks per GPU Processing Cluster. We use the standard data parallel scheduler for non-cooperative kernels and both the standard data parallel and Stream-K streamk schedulers for cooperative ones. All benchmarks are conducted on an NVIDIA H100 SXM GPU.
Our NestedFP16 kernels are optimized using the same design space as the FP16 baseline, ensuring that each GEMM shape is paired with its optimal kernel configuration. We benchmark both kernels across all linear-layer GEMM shapes used in LLaMA 3.1–8B, Mistral Nemo, Phi-4, and Mistral Small, each of which involves four distinct $(N,K)$ shapes. The batch dimension $(M)$, determined by the number of input tokens during inference, is swept from 32 to 2048 in increments of 32, yielding a total of 882 GEMM configurations across 14 unique $(N,K)$ shapes.
Figure 7 compares the performance of the largest $(N,K)$ GEMM configurations across the four models. As shown, our kernel consistently maintains low SIMT overheads across all $M$ values. Compared to the CUTLASS baseline, the average relative overheads across all GEMM configurations of each model are 5.69% (LLaMA 3.1–8B), 6.83% (Mistral Nemo), 6.65% (Phi-4), and 6.71% (Mistral Small). Additional results for other GEMM shapes, as well as comparisons with PyTorch W16A16 kernels, which exhibit similar trends, are provided in Appendix. While certain kernel performance regarding certain $M$ show more variability, the average overhead remains modest, demonstrating the practicality of our approach across diverse shapes in real LLM inference scenario.
Impact of Optimization Techniques. To quantify the impact of successive kernel-level optimizations, we measure latency for a matrix multiplication of shape $M\times 5120\times 32768$. Figure 7 illustrates the latency reduction achieved across three optimization stages. Level 1 introduces a three-stage pipeline that overlaps shared memory transfers with SIMT execution and tensor core operations. Level 2 incorporates SIMT-level FP16 reconstruction optimization. This transition yields a substantial 38.3% reduction in latency relative to Level 1. Level 3 applies additional pipelining and scheduling optimizations resulting in a further 11.0% reduction compared to Level 2. Each stage contributes a measurable improvement, highlighting the cumulative benefits of the proposed optimization strategies.
Figure 7: (a) CUTLASS baseline and NestedFP16 kernel performance comparison. (b) Kernel latency reduction with different levels of optimization strategies.
5.3 End-to-End Performance
We integrated NestedFP into the vLLM framework as a quantization configuration. The system supports per-iteration precision switching, allowing the scheduler to dynamically select FP16 or FP8 based on workload and SLO demands. To ensure fair comparison, both FP16 baselines and NestedFP models were compiled using the vLLM 0.8.3 V1 engine and PyTorch 2.6.0+cu12.4. Experiments were run on a single NVIDIA H100 GPU with fixed request sizes (256 input, 512 output tokens) and batch sizes ($M$) ranging from 32 to 512 to assess throughput scaling. Throughput is reported in tokens per second (tok/s). Figure 8 shows end-to-end performance under FP16 and NestedFP. The FP16 runs use torch.matmul, which dispatches to cuBLAS by default within the vLLM framework. Compared to this baseline, NestedFP16 introduces only minor overhead—4.50% (LLaMA 3.1–8B), 4.51% (Mistral Nemo), 2.69% (Phi-4), and 3.91% (Mistral Small)—generally lower than kernel-level overheads, likely due to amortization across non-GEMM components. In contrast, NestedFP8 achieves clear speedups over NestedFP16: 1.24$\times$, 1.32$\times$, 1.29$\times$, and 1.53$\times$, respectively. Larger models gain more, reflecting the growing cost of matrix multiplication. Overall, NestedFP improves efficiency for large-scale LLMs.
Figure 8: End-to-end performance comparison with four models using FP16, NestedFP16 and NestedFP8.
@@ 6 Conclusion
In this paper, we introduce NestedFP, a precision-aware serving framework designed to better satisfy service-level objectives (SLOs) under dynamic request conditions. We propose a novel approach for decomposing FP16 weights into two 8-bit tensors, enabling one to be used directly as an FP8 representation. Our FP8 format achieves accuracy comparable to that of standard quantized FP8 models. In addition, we develop a custom 16-bit GEMM kernel tailored to this dual-format representation. Experimental results show that our kernel achieves an average performance gap of 6.47% compared to the CUTLASS baseline kernel. Furthermore, by integrating our method into the vLLM framework, we observe up to 1.55$\times$ speedup in FP8 model execution relative to FP16 model execution, while incurring average 3.90% overhead across models when executing in FP16 mode. These results demonstrate that NestedFP successfully bridges the gap between FP16 accuracy and FP8 efficiency, requiring the same memory footprint as a standard FP16 model while supporting both precision modes. By enabling iteration-level precision selection, our approach allows production LLM serving systems to dynamically balance performance and accuracy requirements under varying workload conditions, providing a flexible and efficient solution for real-world deployment scenarios.
@@ Appendix A Evaluation Details
We conducted our evaluation using the Mistral Nemo Base 2407 and Mistral Small 24B Base 2501 models. For accuracy assessment, we employed the Math Verify Metric on the Minerva Math task, the Open LLM Leaderboard configuration (5-shot, multiple choice) for the MMLU Pro task, and a zero-shot evaluation setting for the BBH task. Kernel performance was measured using NVIDIA Nsight Compute. In order to get optimal kernel, we padded activation in the multiple of tile dimension $T_{m}$ as it provides more robust performance. During performance measurement, we flushed all caches using the cache-control option, but did not fix the GPU clock frequency using the clock-control option.
Appendix B Extended Kernel Performance for NestedFP16
Figure 9: Performance comparison between our CUTLASS baseline and NestedFP16 kernel across 14 unique $(N,K)$ GEMM shapes. Each subplot represents a different shape configuration, with batch size $(M)$ on the $x$-axis and latency on the $y$-axis.

### v2

@@ 4.2 FP16 GEMM Kernel with On-the-fly Reconstruction
Figure 5 illustrates the high-level execution flow of the NestedFP FP16 GEMM kernel compared to a standard FP16 GEMM kernel. We assume an NVIDIA GPU environment, where the CUTLASS implementation serves as the standard FP16 GEMM kernel, upon which our kernel is also built. In the standard FP16 GEMM kernel, weights are loaded into shared memory, transferred to registers, and directly consumed by compute units (e.g., tensor cores in NVIDIA GPUs) without any transformation. In contrast, the FP16 GEMM in NestedFP requires a reconstruction step before tensor core execution; weights are stored as two 8-bit tensors, which must be loaded into shared memory, copied to registers, and then reconstructed into full FP16 values. Only after this reconstruction can the data be used by the tensor cores.
Not to make this additional step of reconstruction introduce a bottleneck, we present two key optimizations: (1) merged bitwise operations and (2) a three-stage pipeline.
Efficient FP16 Reconstruction via Merged Bitwise Operations. FP16 reconstruction involves a large number of 8-bit bitwise operations. To reduce the associated overhead, we fuse four 8-bit operations into a single 32-bit instruction, following the approach in any_precision_llm ; fp6_llm . Algorithm 1 illustrates how FP16 reconstruction is performed using merged bitwise operations. Conceptually, the process iterates over the weight tensor, where each iteration processes four elements. In each iteration, four 8-bit values are loaded from the upper and lower tensors ($W_{\text{upper}}$ and $W_{\text{lower}}$) into 32-bit registers u and l, respectively (Line 5–6). Then, the sign bits are extracted from u and stored in another register s (Line 7). Next, the algorithm checks whether rounding has occurred by examining l (Line 8); if rounding is detected, the effect is reversed, thereby recovering the exponent and mantissa parts of the upper 8 bits (Line 9). Then, the original upper 8 bits, u_orig, are fully reconstructed by concatenating them with s (Line 10). Finally, u_orig and l are fused using NVIDIA’s __byte_perm instruction (lines 11–12), completing the reconstruction. Note that all the above operations are simultaneously applied to four 8-bit values through a 32-bit instruction, effectively reducing the total number of bitwise operations required.
@@
Algorithm 1 FP16 Weight Reconstruction
[…]
2WW: reconstructed FP16 weight tensor
3NN: total number of elements in the weight tensor
4for $i=0$ to NN step 44 do
5uint32_t u $W_{\text{upper}}[i:i+4]$
6uint32_t l $W_{\text{lower}}[i:i+4]$
7uint32_t s $u\mathbin{\&}\texttt{0x80808080}$
8uint32_t sub $(l\mathbin{\&}\texttt{0x80808080})\gg 7$
9uint32_t u_orig $(((u-sub)\gg 1)\mathbin{\&}\texttt{0x3f3f3f3f})$
10u_orig $\texttt{u\_orig}\mathbin{|}\texttt{s}$
11W[i:i+2]W[i:i+2] $byte\_perm(\texttt{u\_orig},l,\texttt{0x1504})$
12W[i+2:i+4]W[i+2:i+4] $byte\_perm(\texttt{u\_orig},l,\texttt{0x3726})$
13end for
Three-Stage Pipeline. The CUTLASS FP16 GEMM kernel, upon which the NestedFP FP16 kernel is built, employs a two-stage pipeline to overlap data transfer and computation. Specifically, the two stages correspond to (1) data movement between shared memory and registers and (2) Tensor Core execution, which are overlapped for high throughput. NestedFP extends this design into a three-stage pipeline to further hide reconstruction overheads. In particular, the three stages in NestedFP are: (1) shared-memory-to-register data transfer, (2) FP16 reconstruction, and (3) Tensor Core execution.
@@ 5 Evaluation
Through extensive experiments, we demonstrate that NestedFP is an effective solution for handling fluctuating loads in LLM serving by dynamically switching between FP16 and FP8 modes, as evidenced by the following three key results:
  • The quality of FP8 model extracted from FP16 parameters in the NestedFP data format matches state-of-the-art FP8 quantization results. (Section 5.1)
  • The FP16 mode of NestedFP achieves throughput comparable to baseline FP16 inference using the vanilla CUTLASS kernel, despite performing on-the-fly FP16 reconstruction. (Section 5.2 and Section 5.3)
  • NestedFP significantly improves SLO attainment over FP16-only deployment, without introducing additional memory overhead. (Section 5.4)
5.1 FP8 Model Quality
@@
Methodology. We evaluate the accuracy of the FP8 model in NestedFP against FP8 models quantized using commonly adopted configurations, specifically per-channel weight quantization. For activation quantization, we apply per-token quantization for both NestedFP and the baseline. The evaluation covers six models: Llama 3.1 8B, Mistral Nemo (12B), Phi-4 (14B), Mistral Small (24B), Llama 3.1 70B, and DeepSeek-R1-Distill-Llama-70B deepseek_ref ; nemo_ref ; small_ref ; llama3_technical_report ; phi_ref . For benchmarks, we use three downstream tasks: Minerva Math, MMLU Pro, and BBH minerva ; bbh ; mmlupro . All models are integrated into the vLLM framework, and accuracy is measured using the LM Evaluation Harness lm-evaluation-harness . As noted in Section 4.1, 8.75% of the linear layers in Phi-4 are not applicable to NestedFP and are therefore executed in FP16. More details on the experimental setup are provided in Appendix A.
Results. Table 2 presents the results. Across all models and benchmarks, the FP8 model in NestedFP achieves accuracy comparable to the baseline FP8 model, demonstrating the effectiveness of the new data format of NestedFP.
@@
Llama 3.1 8B Mistral Nemo Phi-4 Mistral Small Llama 3.1 70B DeepSeek-R1- Distill-Llama-70B FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) Minerva Math 17.6 17.2 16.5 16.4 42.9 42.9 34.2 35.2 35.7 34.6 46.0 45.5 MMLU Pro 32.8 33.2 35.3 34.7 52.7 53.3 53.6 53.7 48.1 47.4 49.2 49.4 BBH 38.6 39.0 40.6 40.0 27.1 28.7 50.8 51.1 49.1 49.7 53.2 53.4
Table 2: Accuracy of FP8 models in NestedFP across various downstream tasks, compared to the baseline FP8 model. FP8(B) denotes the baseline model, and FP8(N) denotes NestedFP.
5.2 FP16 GEMM Kernel Performance
Methodology. We evaluate the FP16 GEMM kernel performance of NestedFP against the CUTLASS FP16 GEMM kernel, upon which ours is built, using an NVIDIA H100 GPU. Note that NestedFP does not require any custom kernel for FP8 GEMM; thus, we omit FP8 kernel evaluation. We consider four different weight matrix shapes: (N, K) = (28672, 4096), (28672, 5120), (35840, 5120), (65536, 5120). The $(M)$ dimension, which corresponds to the number of tokens processed in parallel, is varied from 32 to 2048 in increments of 32. For both kernels, we select the best configuration (e.g., tile size) for each GEMM shape through an exhaustive search over the design space. Details of the kernel search space are provided in Appendix D.2.
Results. Figure 6 shows the results. The NestedFP kernel consistently demonstrates performance comparable to the CUTLASS baseline across all GEMM shapes. Specifically, the average relative overheads for each weight matrix dimension are 5.89%, 6.33%, 4.59%, and 6.33%, respectively. These results indicate that the NestedFP kernel effectively minimizes the overhead introduced by on-the-fly FP16 reconstruction. Additional results for other GEMM shapes, which exhibit similar trends, as well as comparisons with cuBLAS FP16 kernels, are provided in Appendix B and Appendix D.1.
Figure 6: CUTLASS baseline and NestedFP kernel performance comparison.
5.3 FP16 Mode End-to-End Inference Performance
Methodology. We evaluate the end-to-end inference throughput of the FP16 mode of NestedFP by integrating it into the vLLM framework on a single NVIDIA H100 GPU. Specifically, we set the input and output sequence lengths to 1024 and 512, respectively, while varying the batch size from 32 to 512. More details on the experimental setup are provided in Appendix C. The results are compared against vanilla vLLM FP16 execution, which relies on PyTorch for invoking GEMM kernels. Again, the FP8-mode throughput of NestedFP is identical to baselines such as vanilla vLLM and is therefore omitted from evaluation.
Results. Figure 7 presents the results. NestedFP achieves FP16-mode throughput highly comparable to the baseline, with only minor overheads—5.04% (LLaMA 3.1–8B), 5.56% (Mistral Nemo), 4.76% (Phi-4), and 4.56% (Mistral Small) on average across batch sizes. The end-to-end throughput degradation is even lower than the kernel-level overheads reported in Section 5.2, since non-GEMM components also contribute to total execution time. Additional end-to-end throughput results are provided in Appendix C.
Figure 7: FP16 inference throughput comparison of four models between vanilla vLLM and NestedFP.
5.4 Impact on SLO Attainment
Methodology. In this section, we present a case study demonstrating how NestedFP can improve SLO attainment. We implement a simple load-aware mode switching policy for NestedFP that, in each iteration, selects FP8 mode when the number of tokens to process exceeds a predefined threshold (1024 in this experiment); otherwise, it selects FP16 mode. The evaluation is conducted using Llama 3.1–70B on a cluster equipped with four NVIDIA H100 GPUs. As input, we use 1,000 requests sampled from an Azure LLM inference trace while varying the request rate. We measure p90 TTFT, p90 TPOT, and SLO attainment under two criteria: loose and tight. The loose SLO is defined as $\leq 10\times$ the average single-request latency, and the tight SLO is defined as $\leq 5\times$ the average single-request latency, following the methodology of a previous study sola_llm .
Results. Table 3 shows the results. NestedFP consistently outperforms the FP16-only baseline across all request rates in terms of TTFT, TPOT, and SLO attainment, with particularly pronounced improvements under high request rates. Note that this experiment employs a simple, straightforward switching policy, leaving substantial room for further optimization by integrating NestedFP with more sophisticated policies—such as those considering KV-cache utilization or incorporating load prediction mechanisms. Nevertheless, these results confirm that NestedFP provides an efficient dual-precision mechanism that can significantly enhance SLO attainment in LLM serving systems.
Request Rate Method Tight SLO (%) Loose SLO (%) p90 TTFT (s) p90 TPOT (ms)
7.47 req/s FP16 0.7 14.1 9.9893 162.2
NestedFP 72.4 90.0 1.9518 118.9
5.03 req/s FP16 71.6 93.6 1.6449 151.3
NestedFP 94.6 99.9 0.9679 86.8
3.79 req/s FP16 83.1 96.9 1.2404 109.2
NestedFP 96.6 100 0.8168 60.1
1.99 req/s FP16 90.4 99.4 0.8915 64.9
NestedFP 98.4 100 0.6116 47.7
Table 3: SLO attainment and p90 TTFT/TPOT of Llama-3.1-70B on 4×H100 under varying loads.
@@ 6 Conclusion
We introduce NestedFP, a memory-efficient framework that supports both FP16 and FP8 precisions for LLM inference in a memory-efficient way. NestedFP introduces a new data format that enables direct extraction of FP8 weights from FP16 weights without additional memory overhead. NestedFP also incorporates a custom GEMM kernel optimized for this format to ensure efficient computation. As a result, NestedFP enables dynamic switching between FP16 and FP8 modes with the memory footprint of FP16, without noticeable degradation in either accuracy or inference throughput, providing an effective means to improve SLO attainment in LLM serving.
7 Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea Government (MSIT) (RS-2024-00340008) and Institute of Information & Communications Technology Planning & Evaluation (IITP) under the artificial intelligence semiconductor support program (IITP-2023-RS-2023-00256081), funded by the Korea Government (MSIT).
@@ Appendix A Evaluation Details
We conducted our evaluation using the Mistral Nemo Base 2407 and Mistral Small 24B Base 2501 models. For accuracy assessment, we employed the Math Verify Metric on the Minerva Math task, the Open LLM Leaderboard configuration (5-shot, multiple choice) for the MMLU Pro task, and a zero-shot evaluation setting for the BBH task. Kernel performance was measured using NVIDIA Nsight Compute. In order to get optimal kernel, we padded activation in the multiple of tile dimension $T_{m}$. During performance measurement, we flushed all caches using the cache-control option, but did not fix the GPU clock frequency using the clock-control option.
Appendix B Extended Kernel Performance for NestedFP
Figure 8: Performance comparison between our CUTLASS baseline and NestedFP kernel across 14 unique $(N,K)$ GEMM shapes. Each subplot represents a different shape configuration, with batch size $(M)$ on the $x$-axis and latency on the $y$-axis.